# Detecting and Filtering Unsafe Training Data via Data Attribution

### v1

@@ Abstract
Large language models (LLMs) are vulnerable to unsafe training data that even small amounts of unsafe data can lead to harmful model behaviors. Detecting and filtering such unsafe training data is essential for trustworthy model development. Current state-of-the-art (SOTA) approaches typically rely on training moderation classifiers which requires significant computational overhead and are limited to predefined taxonomies, making them less adaptable to evolving safety concerns. Moreover, these classifiers lack insight into the training process, limiting their effectiveness in filtering unsafe data. To address these limitations, we propose DABUF, leveraging data attribution to detect and filter unsafe training data by attributing harmful model outputs to influential training data points. DABUF enables flexible identification of various unsafe data types without predefined taxonomies. However, in practice, model outputs can be complex with combined safe linguistic features and unsafe content, leading to reduced attribution accuracy. In such cases, DABUF will integrate moderation classifiers to identify a minimal subset of unsafe training data for targeted attribution (such as jailbreak). When model outputs are relatively straightforward, DABUF uses model outputs directly as the attribution targets. We evaluate the performance on two different tasks: in filtering jailbreaking training data and in identifying and mitigating gender bias. DABUF outperforms SOTA approaches by up to 7.5% in detection AUPRC in jailbreaking scenarios, and 44.1% in detecting gender bias. Moreover, retraining on DABUF-filtered data leads to higher model safety across experiments, underscoring its versatility in addressing a broad spectrum of unsafe data issues.
Detecting and Filtering Unsafe Training Data via Data Attribution
@@ 1 Introduction
Large Language Models (LLMs) are known to exhibit various unsafe behaviors, including toxicity, stereotyping, privacy leaks, and ethical violations (Wang et al., 2024a). A primary source of these issues is unsafe training data (Jiang et al., 2024; Chen et al., 2024). For instance, inherent biases or toxic content in a dataset can lead to harmful responses (Jiang et al., 2024; Ouyang et al., 2022), while deliberate attacks, such as adversarial prompts or injected backdoors, can be used to bypass safety alignments (Chen et al., 2024; Zou et al., 2023; Li et al., 2024b). Consequently, identifying and removing these unsafe training instances is critical for mitigating risks and building safer LLMs.
@@
Existing methods for detecting and filtering unsafe training data typically rely on moderation classifiers. Online API tools, such as the Perspective API  [^1] https://www.perspectiveapi.com/ and OpenAI’s Moderation API (Markov et al., 2023), focus on certain predefined toxicity taxonomies, but struggle to generalize to nuanced and emerging unsafe artifacts beyond these predefined taxonomies Weber et al. (2025). Fine-tuned detection models, including Llama-Guard-3-8B (Llama Team, 2024) and Wildguard (Han et al., 2024), require significant time and resources for additional data collection and training. Moreover, these moderation classifiers primarily detect semantically unsafe training data without considering the influence of each data point on model training, resulting in suboptimal filtering effectiveness for enhancing model safety.
In this work, we introduce Data-Attribution-Based Unsafe Training Data Detection and Filtering (DABUF), a method that leverages data attribution techniques to enhance unsafe data detection and filtering in LLMs. Data attribution is a family of methods that quantify the influence of individual training data points on specific model outputs (Koh and Liang, 2020; Pruthi et al., 2020). Our central hypothesis is that unsafe training data exerts greater influence on unsafe outputs; thus, attributing these outputs back to their influential training instances can reveal which data points are responsible. These methods do not require additional training data and can be applied flexibly to diverse types of unsafe model outputs.
However, applying data attribution naively—by directly attributing unsafe model generations to their training data—is ineffective in many cases. LLM outputs, particularly in long-form generations such as jailbreaking attacks, are influenced by a mix of benign and unsafe training data. Since model generations include common linguistic structures (e.g., stop words, neutral phrases) alongside unsafe content, direct attribution to the entire sequence leads to a noisy attribution signal, reducing the precision of unsafe data identification.
To address this challenge, our method introduces a targeted filtering mechanism. Specifically, for long-form outputs, such as those found in jailbreaking scenarios, we first use moderation classifiers to identify a small subset of clearly unsafe training data. We then use this subset as the attribution target, allowing us to refine the attribution process and isolate the most influential unsafe training instances. In contrast, for shorter model outputs, such as those in gender bias scenarios, where the influence of training data is more direct and less noisy, standard data attribution techniques are sufficient without additional moderation filtering.
We validate our approach through experiments in two distinct setups. In jailbreaking scenarios involving adversarial prompts that lead to noisy or long unsafe outputs, we apply the proposed DABUF with moderation classifiers for initial unsafe data identification. Conversely, in the gender bias scenario—where outputs are relatively concise—we directly apply DABUF on the model generation. Experimental results show that our approach achieves superior detection performance across different model architectures in jailbreaking scenarios and deliver improved safety when models are retrained with filtered data, outperforming state-of-the-art detection methods. Furthermore, our method generalizes effectively to gender bias scenarios, highlighting its versatility.
@@ 2.2 Unsafe Training Data Detection in LLMs
Beyond content moderation classifiers, some recent studies have leveraged the internal structures of models for unsafe data detection. For example, GradSafe (Xie et al., 2024) utilizes gradient similarity with respect to safety-critical parameters to identify unsafe data, while BEBC (Zheng et al., 2024) employs fine-tuned BERT embeddings for content moderation.
Unlike online moderation tools and LLM-based classifiers, our methods eliminate the need for data curation and additional training. Furthermore, by operating independently of predefined safety taxonomies, our approach demonstrates greater flexibility and can effectively handle a wider range of safety-related scenarios. In contrast to other non-classifier approaches, such as GradSafe, our methods address the problem through the lens of data attribution. This perspective enhances detection performance by incorporating unsafe model behavior and capturing the relationship between training data and model outputs, ultimately fostering safer model developments.
2.3 Data Attribution for LLMs
Data attribution methods aim to quantify the impact of individual training samples on a model’s predictions for specific test cases (Koh and Liang, 2020). These methods have the potential to detect unsafe training data, as such data are likely to exert a disproportionate influence on unsafe model outputs, making them distinguishable from the broader benign dataset. Recently a variety of data attribution methods have been proposed to estimate the influence of training data in the context of LLMs. These include gradient-based methods (Xia et al., 2024; Kwon et al., 2024), simulator-based methods (Guu et al., 2023; Chai et al., 2024) and game-theoretic methods (Wang et al., 2024b, c). Estimated influence scores have been utilized for tasks such as identifying mislabeled data (Pruthi et al., 2020), understanding memorization (Feldman and Zhang, 2020), and data valuation (Choe et al., 2024). While data attribution methods have various applications in LLMs, they are computationally intensive, limiting their applicability to larger models. Despite recent advancements in efficient influence estimation methods Kwon et al. (2024), the computational burden remains a challenge. Gradient-similarity-based approaches, as highlighted in previous works Xia et al. (2024); Pruthi et al. (2020), offers an efficient solution, making it better suited for scaling to LLMs.
3 Data-Attribution-Based Unsafe Training Data Detection and Filtering
Our proposed method consists of two phases: detection and filtering. The primary technical challenge arises in the detection phase, where we identify unsafe data points in the training dataset that contribute to unsafe model behaviors. In the filtering phase, we mitigate these behaviors by removing the data points most likely to be unsafe.
3.1 Unsafe Training Data Detection
We first describe the problem setup of unsafe training data detection. Consider a training dataset with a mixture of benign and unsafe data:
@@ 3.1 Unsafe Training Data Detection
where $\mathcal{D}_{\text{benign}}$ refers to the benign dataset that is safe to train on while $\mathcal{D}_{\text{unsafe}}$ is the unsafe training dataset that could lead to unsafe model behaviors. In addition, we assume access to a (small) target dataset $\mathcal{D}_{\text{target}}$ that consists of unsafe model outputs or examples of the unsafe training data. In practice, these examples may come from user reports or manual inspection of a small portion of training data.
The goal of unsafe training data detection is to retrieve $\mathcal{D}_{\text{unsafe}}$ from the entire $\mathcal{D}_{\text{train}}$ with high precision and recall, possibly using the information from the target set $\mathcal{D}_{\text{target}}$. A high-quality detection method will help us obtain a cleaner training dataset without overly removing safe training data in the filtering phase.
Data-Attribution-Based Detection
We propose to detect the unsafe training data by measuring the influence of each training data point $z\in\mathcal{D}_{\text{train}}$ on the likelihood of the model generating the examples in the target dataset $\mathcal{D}_{\text{target}}$. Intuitively, since $\mathcal{D}_{\text{target}}$ consists of unsafe examples, a training data point with higher influence is more likely to be unsafe. Formally, we denote the influences as
$\text{Inf}(z,\mathcal{D}_{\text{target}}),\quad z\in\mathcal{D}_{\text{train}}.$
In this work we use gradient similarity (Pruthi et al., 2020) to efficiently estimate training data’s influence on model generations, which is a scalable method that has been widely used in data attribution for LLMs (Xia et al., 2024). Consider a model parameterized by $\theta$, and denote the negative log-likelihood as $\ell(\cdot;\theta)$. The influence of a training data point $z\in\mathcal{D}_{\text{train}}$ on the target dataset $\mathcal{D}_{\text{target}}$ is defined as following:
$\mathrm{Inf}(z,\mathcal{D}_{\text{target}}):=\eta\cos(\nabla\ell(\mathcal{D}_{\text{target}};\theta),\nabla\ell(z;\theta)),$
where $\eta$ is the average learning rate, $\nabla\ell(\mathcal{D}_{\text{target}};\theta)$ is the gradient of the negative log-likelihood with respect to $\theta$ evaluated on the target set $\mathcal{D}_{\text{target}}$, and $\nabla\ell(z;\theta)$ is the gradient evaluated on the training point $z$. To improve computational efficiency in practice, we follow Xia et al. (2024) to reduce the gradient dimension to $d=8192$ via random projection, and adopt optimizer-aware training gradients.
Data Attribution for LLM Outputs
In the context of LLMs, the model output is a sequence of tokens. Specifically, for each example $x\in\mathcal{D}_{\text{target}}$, it can be represented as $x=(p,r)$, where $p$ is the input prompt and $r$ is the output response, both of which are a sequence of tokens. Existing literature (Xia et al., 2024) typically defines $\nabla\ell(\mathcal{D}_{\text{target}};\theta)$ as
$\nabla\ell(\mathcal{D}_{\text{target}};\theta)=\sum_{x\in\mathcal{D}_{\text{target}}}\nabla\log p(r|p;\theta),$
where one can further expand the conditional probability $p(r|p;\theta)$ defined by an autoregressive LLM as following:
$\nabla\ell(r;\theta,p)=\sum_{i=1}^{|r|}\nabla\ell(r_{i}|p,r_{<i};\theta).$
However, we find that naively applying the data attribution method defined above to long-form unsafe outputs often yields suboptimal performance for identifying unsafe training data. When a response $r$ contains many benign tokens and only a few segments of genuinely unsafe content, the attribution signal becomes diluted by the larger volume of neutral or benign tokens. In particular, the tokens directly associated with unsafe content–henceforth referred to as unsafe tokens–carry disproportionately stronger gradients, indicating their direct link to harmful outputs. Yet, when the model’s overall attribution signal is aggregated over all tokens, these critical unsafe signals become overwhelmed by the contributions of benign tokens, resulting in noisy and less precise detection of unsafe training data points. This imbalance is especially problematic in long-form scenarios like jailbreaking attacks, where substantial portions of the model response may be benign filler text interspersed with targeted unsafe content. Consequently, the naive approach of attributing every token in the entire generation fails to isolate the truly influential unsafe training instances. Please refer to Appendix B.1 for more detailed empirical evidence of this observation.
Leveraging Externally-Identified Unsafe Data for Effective Attribution
To address the aforementioned issue in scenarios where the model outputs are long, we propose to take ground-truth labels (instead of the model outputs) from a small, externally identified subset of unsafe training data as the target dataset to attribute. For example, in the jailbreaking scenario, we first an LLM-based classifier (Llama-3-Guard-8B (Llama Team, 2024)) to screen the entire training dataset and obtain a small candidate set of unsafe training data $\mathcal{D}_{\text{cand}}$. Because such classifiers can have high false-positive rates, we further perform human annotation on $\mathcal{D}_{\text{cand}}$ to filter out benign data points, resulting in a smaller, verified unsafe subset $\mathcal{D}_{\text{identified}}$. The target set is then set as $\mathcal{D}_{\text{target}}=\mathcal{D}_{\text{identified}}$. In all of our experiments, $|\mathcal{D}_{\text{cand}}|$ is well below 200–manageable for human inspection–whereas the full training set exceeds 40,000 samples.

### v2

@@ Abstract
Large language models (LLMs) are highly sensitive to even small amounts of unsafe training data, making effective detection and filtering essential for trustworthy model development. Current state-of-the-art (SOTA) detection approaches primarily rely on moderation classifiers, which require significant computation overhead for training and are limited to predefined taxonomies. In this work, we explore data attribution approaches that measure the similarity between individual training samples and a small set of unsafe target examples, based on data representations such as hidden states or gradients. We identify a key limitation in existing methods: unsafe target texts contain both critical tokens that make them unsafe and neutral tokens (e.g., stop words or benign facts) that are necessary to form fluent language, and the latter of which makes the overall representations “noisy” for the purpose of detecting unsafe training data. To address this challenge, we propose Denoised Representation Attribution (DRA), a novel representation-based data attribution approach that denoises training and target representations for unsafe data detection. Across tasks of filtering jailbreaks and detecting gender bias, the proposed approach leads to significant improvement for data attribution methods, outperforming SOTA methods that are mostly based on moderation classifiers.
Detecting and Filtering Unsafe Training Data via Data Attribution with Denoised Representation
@@ 1 Introduction
Large Language Models (LLMs) are known to exhibit various unsafe behaviors, including toxicity, stereotyping, privacy leaks, and ethical violations (Wang et al., 2024a). A primary source of these issues is unsafe training data (Jiang et al., 2024; Chen et al., 2024). For instance, inherent biases or toxic content in training data can lead to harmful responses (Jiang et al., 2024; Ouyang et al., 2022), while deliberate training-time attacks, such as adversarial prompts or injected backdoors, can be used to bypass safety alignments (Chen et al., 2024; Zou et al., 2023; Li et al., 2024b). Consequently, identifying and removing these unsafe training instances is critical for mitigating risks and building safer LLMs.
@@
Existing methods for detecting and filtering unsafe training data typically rely on moderation classifiers. Online API tools, such as the Perspective API  [^1] https://www.perspectiveapi.com/ and OpenAI’s Moderation API (Markov et al., 2023), focus on certain predefined toxicity taxonomies, struggling to generalize to nuanced and emerging unsafe artifacts beyond these predefined taxonomies Weber et al. (2025). Fine-tuned detection models, including Llama-Guard-3-8B (Llama Team, 2024) and Wildguard (Han et al., 2024), require significant time and resources for additional data collection and training. Moreover, these moderation classifiers primarily flag stylistically unsafe data in a model-agnostic way, without accounting for how individual training samples actually influence the model. As a result, they fail to capture the true training impact of unsafe data and are less effective for enhancing model safety through retraining.
Another line of research focuses on filtering training data by measuring data representation similarity with a set of unsafe model responses. These methods are appealing because they do not require additional labeled data and can be flexibly applied across diverse forms of unsafe model outputs. The core hypothesis is that unsafe training data exhibit higher similarity to unsafe model outputs. Prior work has explored model hidden states, for example, LARF Li et al. (2025) and RepSim Pezeshkpour et al. (2021), as well as gradient-based similarity (Xia et al., 2024). These methods are widely adopted in the data attribution literature (Deng et al., 2025) to quantify the influence of training data on target data (Pruthi et al., 2020).
However, existing representation-based methods typically measure similarity between training samples and target unsafe examples using representations averaged over all tokens in each sample. This naive averaging introduces a key limitation: not all token representations contribute equally to detecting unsafe data. Unsafe texts often contain only a few critical tokens that account for the unsafe semantics while the majority of tokens (e.g., stop words or neutral phrases) are benign and contribute to linguistic fluency or neutral factual information. As a result, the averaged representations are “noisy”, obscuring the true signal needed to identify unsafe training data.
To address this limitation, we propose Denoised Representation Attribution (DRA), a novel representation-based data attribution method for unsafe training data detection and filtering. DRA denoises the training and target representations in two steps. First, motivated by the optimal linear detector in signal detection theory (Stanislaw and Todorov, 1999), DRA centers and whitens the data representations so spurious high-variance dimensions often irrelevant to the unsafeness are downweighted. Second, DRA applies a simple leave-target-out surrogate to select most informative hidden dimensions of the representations for unsafe training data detection. This two-step procedure leads to denoised representations that are more effective to detect unsafe training samples.
We validate our proposed approach through experiments on both jailbreaking and gender bias scenarios. Experimental results demonstrate that our method consistently outperforms state-of-the-art detection approaches across different model architectures, increasing average detection AUPRC by up to $63.3\%$. Moreover, when retraining models on data filtered by our method, we decrease average ASR (Attack Success Rate) by up to $39.9\%$ , confirming the effectiveness of our approach in mitigating unsafe training data.
@@ 2.2 Unsafe Training Data Detection in LLMs
Beyond safety classifiers, other approaches rely on representation-based similarity for unsafe data detection. For example, BEBC (Zheng et al., 2024) employs fine-tuned BERT embeddings for content moderation, while gradient-based methods such as GradSafe (Xie et al., 2024) estimate similarity with respect to safety-critical parameters. More broadly, gradient-similarity methods can be seen as a form of data attribution (Koh and Liang, 2020), which quantify the influence of individual training samples on model outputs. These attribution-based perspectives are particularly useful for detecting unsafe data, as such data often exert disproportionate influence on unsafe behaviors, making them distinguishable from the broader benign dataset. Although influence estimation can be computationally intensive, recent gradient-similarity approaches (Xia et al., 2024; Kwon et al., 2024) provide more efficient alternatives, making them better suited for scaling to large models.
However, prior work such as GradSafe and BEBC primarily focus on prompt filtering, whereas in training data detection the training label is also important, since it drives model behavior during training. Naively extending these methods to the training data setting yields suboptimal results, as noise in model responses manifests as degraded detection performance (Jiao et al., 2025). In this context, our methods enable efficient and principled denoising and dimension selection techniques for representation-based detection. Our approach enhances detection performance and fosters safer model developments.
3 Unsafe Training Data Detection and Filtering
In this section, we formally define the task of unsafe training data detection, introduce a framework for quantifying the sensitivity of representation-based detection methods, and present our proposed Denoised Representation Attribution (DRA) approach for improving detection effectiveness. We also describe how this detection method can be used to filter unsafe data for safer downstream model training.
3.1 Problem Setup: Representation-Based Unsafe Training Data Detection
Consider a training dataset with a mixture of benign and unsafe data:
@@ 3.1 Problem Setup: Representation-Based Unsafe Training Data Detection
where $\mathcal{D}_{\text{benign}}$ refers to the benign dataset that is safe to train on while $\mathcal{D}_{\text{unsafe}}$ is the unsafe training dataset that could lead to unsafe model behaviors. In real world scenarios, the training dataset often consists of only a small portion of unsafe data, and the membership of $\mathcal{D}_{\text{benign}}$ and $\mathcal{D}_{\text{unsafe}}$ is unknown. The goal of unsafe training data detection is to identify a subset of training data points in $\mathcal{D}_{\text{train}}$ that are mostly likely to be unsafe.
In addition, we assume access to a (small) target dataset $\mathcal{D}_{\text{target}}$ that consists of unsafe model outputs. In practice, these examples may be flagged by users or collected via red teaming.
Representation-based detection.
For any $z\in\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{target}}$, let $h$ be the mapping from the data point $z$ to its representation $h(z)\in\mathbb{R}^{m}$ (e.g., hidden states or gradients). Then we define $B=\{h(z)\mid z\in\mathcal{D}_{\text{benign}}\},U=\{h(z)\mid z\in\mathcal{D}_{\text{unsafe}}\},$ and $Q=\{h(z)\mid z\in\mathcal{D}_{\text{target}}\},$ which are respectively the set of data representations for $\mathcal{D}_{\text{benign}},\mathcal{D}_{\text{unsafe}},$ and $\mathcal{D}_{\text{target}}.$ The set of data representations for $\mathcal{D}_{\text{train}}$ is then $X=B\cup U$.
A common paradigm in prior work Xie et al. (2024); Xia et al. (2024); Li et al. (2025) is to assign each training representation $x\in X$ ($x\in\mathbb{R}^{m}$) a detection score based on its similarity to an aggregated target representation $\mu_{Q}=\mathrm{Agg}(Q)\in\mathbb{R}^{m}$ (e.g., $\mathrm{Agg}(Q)=\frac{1}{|Q|}\sum_{q\in Q}q$). Training samples with the highest detection scores are then flagged as unsafe. Formally, the score is typically defined as the inner product
$s_{\mu_{Q}}(x)\;=\;\langle x,\mu_{Q}\rangle.$
In LLMs, both the training representation $x$ and the aggregated target representation $\mu_{Q}$ encode rich semantic information, much of which is irrelevant to the unsafeness of the samples. For the purpose of detecting unsafe training data, both representations are therefore noisy.
Goal.
Our study aims to quantify the sensitivity of the detection score $s_{\mu_{Q}}$, and develop better representation-based detection measures.
3.2 Sensitivity of the Detection Score
We frame the problem of detecting unsafe training data in the language of signal detection theory Stanislaw and Todorov (1999): the unsafe training data act as the signal to be distinguished from a much larger background of benign data constituting the noise. The sensitivity of such detection can be quantified by the classical discriminability index, $d^{\prime}=\frac{\mu_{\text{signal}}-\mu_{\text{noise}}}{\sigma_{\text{noise}}}$, which depends on the means ($\mu_{\cdot}$) and standard deviations ($\sigma_{\cdot}$) of a given detection measure for the signal and noise distributions, and reflects how well the two distributions can be separated.
By applying this formulation to the detection score $s_{\mu_{Q}}$ in our problem, we obtain the following sensitivity measure for detecting signal samples $U$ from noise samples XX [^2] Strictly speaking, the noise samples are $B$. Under the assumption that $|U|$ is much smaller than $|B|$, the empirical distributions of $X$ and $B$ will be close. using $s_{\mu_{Q}}$:
$d^{\prime}(s_{\mu_{Q}})\;=\;$ $\frac{\mathbb{E}_{u\sim U}[s_{\mu_{Q}}(u)]-\mathbb{E}_{x\sim X}[s_{\mu_{Q}}(x)]}{\sqrt{\mathrm{Var}_{x\sim X}[s_{\mu_{Q}}(x)]}}$
$=\;$ $\frac{\langle\mu_{U}-\mu_{X},\mu_{Q}\rangle}{\sqrt{\mu_{Q}^{\top}\Sigma_{X}\mu_{Q}}},$ (1)
where $\mu_{U}$ and $\mu_{X}$ are the mean of $U$ and $X$, and $\Sigma_{X}=\mathrm{Cov}(X)\;\in\;\mathbb{R}^{m\times m}$ is the covariance of $X$. It is worth noting that under certain technical assumptions, the discriminability index $d^{\prime}$ is monotonic with respect to AUROC and AUPRC Stanislaw and Todorov (1999). Thus $d^{\prime}$ can be viewed as a theoretically grounded surrogate of these widely used evaluation metrics for unsafe data detection, while additionally admitting a natural decomposition that enables fine-grained analysis of the representation-based detection.
3.3 Improving the Detection Score by Denoising Representations
Given the discriminability index $d^{\prime}$, our goal is to develop a representation-based detection score $s$ that achieves a higher $d^{\prime}(s)$.
Optimal linear detection score.
We consider the family of linear detection scores, $s_{w}(x)\;=\;\langle x,w\rangle,$ for any linear weights $w\in\mathbb{R}^{m}$. There is a classical result Poor (2013) in signal detection theory about the optimal linear detection score as stated below.
Lemma 3.1 (Optimal Linear Detection Score)
For any $w\in\mathbb{R}^{m},$ denoting $\Delta=\mu_{U}-\mu_{X}$, then
$d^{\prime}(s_{w})=\frac{\langle\Delta,w\rangle}{\sqrt{w^{\top}\Sigma_{X}w}}$ $=\frac{\langle\Sigma_{X}^{-1/2}\Delta,\Sigma_{X}^{1/2}w\rangle}{\|\Sigma_{X}^{1/2}w\|}$
$\leq\|\Sigma_{X}^{-1/2}\Delta\|,$
with equality achieved for optimal $w^{*}\propto\Sigma_{X}^{-1}\Delta.$
However, note that this theoretically optimal linear weight $w^{*}$ is infeasible to compute, as it depends on $\mu_{U}$, which is unknown due to the unavailability of the unsafe dataset $\mathcal{D}_{\text{unsafe}}$.
Plug-in detection score.
In practice, under the hypothesis that the mean $\mu_{Q}$ from the target dataset aligns well with $\mu_{U}$, we propose to set $\hat{w}=\Sigma_{X}^{-1}(\mu_{Q}-\mu_{X})$, which gives us a plug-in detection score $s_{\hat{w}}$ as an estimate of the optimal linear detection score $s_{w^{*}}$.
Intuitively, compared to the raw score $s_{\mu_{Q}}(x)=\langle x,\mu_{Q}\rangle$, the plug-in score
$s_{\hat{w}}(x)$ $=\big\langle x,\Sigma_{X}^{-1}(\mu_{Q}-\mu_{X})\big\rangle$
$=\big\langle\Sigma_{X}^{-1/2}(x-\mu_{X}),\Sigma_{X}^{-1/2}(\mu_{Q}-\mu_{X})\big\rangle+c,$
is a Mahalanobis correlation (up to a constant $c$ independent of $x$): it first centers both train and target representations by the background mean $\mu_{X}$ to remove global content common to almost all training points, then whitens by $\Sigma_{X}^{-1/2}$ to suppress the high-variance dimensions of the data representation, since the dominant dimensions could correspond to topics or styles instead of the unsafeness.
Dimension selection via leave-target-out discriminability index.
For LLMs, the data representation dimension size $m$ can be very large. Even with the denoising step described above, the representations could still contain too much irrelevant information. Since dimension selection is essentially a subset selection problem, the task becomes challenging when $m$ is large. We further propose to conduct a dimension selection to select the most informative hidden dimensions in the representation for unsafe training data detection.
Specifically, given the eigen decomposition of $\Sigma_{X}=V\Lambda V^{\top}$, we will select a subset of eigenvectors $V_{I}\in\mathbb{R}^{m\times|I|}$ indexed by indices $I\subseteq\{1,2,\ldots,m\}$. Let $\Pi_{I}=V_{I}V_{I}^{\top}$ be the projection matrix for the space spanned by $V_{I}$. We define a detection score restricted to the selected dimensions $I$ as following:
$s_{I}(x)=\big\langle\Pi_{I}\,\Sigma_{X}^{-1/2}(x-\mu_{X}),\;\Pi_{I}\,\Sigma_{X}^{-1/2}(\mu_{Q}-\mu_{X})\big\rangle.$
A natural criterion for selecting the dimensions $I$ is to maximize the discriminability index applied to $s_{I}$:
$d^{\prime}(s_{I})=\frac{\mathbb{E}_{u\sim U}[s_{I}(u)]-\mathbb{E}_{x\sim X}[s_{I}(x)]}{\sqrt{\mathrm{Var}_{x\sim X}[s_{I}(x)]}}.$
However, since $d^{\prime}(s_{I})$ depends on $U$, it is infeasible to compute. Again, we utilize the trick of approximating $U$ from the ground truth unsafe training dataset by $Q$ from the target dataset, resulting in a leave-target-out (LTO) discriminability index:
$d^{\prime}_{\text{LTO}}(s_{I})=\frac{\mathbb{E}_{q\sim Q}[s_{I}(q)]-\mathbb{E}_{x\sim X}[s_{I}(x)]}{\sqrt{\mathrm{Var}_{x\sim X}[s_{I}(x)]}}.$
We name it “leave-target-out” as $d^{\prime}_{\text{LTO}}$ can be viewed as measuring the sensitivity of a hypothetical detection task that aims to identify $\mathcal{D}_{\text{target}}$ out of $\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{target}}$.
The proposed DRA detection score.
All together, the proposed Denoised Representation Attribution (DRA) method is a detection score defined as $s_{I^{*}}(\cdot)$, where $I^{*}=\arg\max_{|I|=r}d^{\prime}_{\text{LTO}}(s_{I})$ for some given dimension size $r$ (we report results for $r=30$ across experiments).
In practice, instead of directly selecting the dimensions from all $m$ eigenvectors, we first reduce the candidate dimensions to the $m^{\prime}<m$ top eigenvectors, and then optimize the indices $I\subseteq\{1,2,\ldots,m^{\prime}\}$.